# Traversability-Aware Legged Navigation by Learning from Real-World Visual Data

### v1

@@ V-A4 Policy Representation and Training Details
The depth-based planner $\pi^{p}_{\text{depth}}$ is a deep neural network with two components: a depth image encoder $\phi_{\text{depth}}$ and a base MLP. We train the depth-based planner $\pi^{p}_{\text{depth}}$ using PPO in Isaac Gym, where random obstacles (cylinders, boxes) and simulated noise are added to the depth images.
@@ V-C Online Learning for Traversability-Aware Navigation
After collecting the offline replay buffer, we begin training the final policy, the traversability-aware planner $\pi^{p}_{\text{rgbd}}$ that fuses depth and color images, and robot state feedback. We use RLPD [21], which allows us to learn from both online data and offline data collected beforehand. This training is from scratch and is directly deployed on the robot hardware while exploring different terrains during navigation.
We retain the depth and color image encoders, $\phi_{\text{depth}}$ and $\phi_{\text{rgb}}$, and only train the base MLP online to avoid the high dimensionality of raw image data and accelerate real-world training. The training objective is to enable the base MLP to properly utilize the latent embeddings from the depth and color images, along with the robot feedback, to avoid terrains with high traversability cost and obstacles while navigating to the goal.
The elements in the POMDP (such as observation, reward, and action) during this stage remain the same as those in the offline data collection stage, as described in Sec. V-B. The key difference is that we now start updating the planner policy (which produces the action $\mathbf{a}^{p}_{t}$) using the data collected with the new policy. We adopt DroQ algorithm [27] (modified SAC [22] with dropout in Q-value network), augmented it with RLPD. During each policy update, the sampled batch consists of two parts: half of the data is sampled from the offline dataset $\mathcal{D}_{\text{offline}}$ previously collected, and the other half is sampled from the online data $\mathcal{D}_{\text{online}}$ produced by the current policy. The update rule follows the standard SAC algorithm.
@@ VI-B Estimating Traversability
As shown in Fig. 5, our method, which includes the value estimate $T_{\text{value}}$ to evaluate overall control performance, performs significantly better than using the velocity tracking error alone ($T$ from Tracking), resulting in a lower total traversability cost over the trajectory (i.e. in sparse obstacles: $2.23\pm 0.52$ s) compared with $T$ from Tracking ($3.14\pm 0.66$ s), which uses the velocity tracking error alone. The results indicate that relying solely on heuristic-based metrics, such as velocity tracking error, fails to account for other factors that may not affect tracking performance but cause other degradations, such as foot slippage or body instability. These factors can be considered by the proposed estimator based on the controller’s value function.
@@ VI-D Different Sources of Offline Data
However, if we only rely on the data collected by the human demonstrators, we also find there is performance degradation, especially for the goal-reaching objective (Ours $0.27\pm 0.06$ m/s versus Human Demo. Only $0.23\pm 0.03$ m/s). This is because human demonstrators may prefer the strategy that favors obstacle avoidance and lower traversability cost, but could be hard to consider minimizing traveling time. The data collected by the depth-based planner can overcome this problem as it explicitly handles the goal vector information with a well-trained planning policy.
@@ VI-E Learning from Both Offline and Online Data
Compared to the Vanilla SAC method, which does not use any pre-collected dataset, we find that incorporating an offline replay buffer, as in our proposed method, significantly improves overall training performance. This advantage is particularly evident in our case, where we are addressing multi-modal learning from high-dimensional visual data. The pre-collected offline replay buffer, $\mathcal{D}_{\text{offline}}$, provides a high-quality dataset that helps to warm up the learning process, enhance the sampling efficiency for online learning, and broaden the coverage of the distribution of collected trajectories.
We also compared our method with another approach that incorporates a pre-collected dataset during learning, but without using reinforcement learning: Behavior Cloning (BC). As shown in Fig. 8, BC demonstrates worse performance because it is more susceptible to suboptimal demonstrations and fails to explore behaviors that could be more optimal than those demonstrated. In contrast, our proposed method is less sensitive to the quality of the demonstrations because the transition pairs collected by our method include a reward term. After online exploration and learning, the agent optimizes for accumulated rewards and surpasses the performance of suboptimal demonstrations.
@@ VII-B Traversability Estimation in the Wild
As shown in Fig. 9, the estimation results accurately reflect changes in terrain. The final result of the scalar traversability cost $T$ significantly increases when the robot traverses cobblestone areas and decreases on nominal concrete roads, correctly reflecting the difficulty of traversing through these regions. For example, cobblestone areas are challenging for the robot to traverse and introduce obvious oscillation on the robot base. The results of this estimatior are used as part of the reward $r^{p}_{t}$ in Eq. (3) to train the planner $\pi_{\text{rgbd}}^{p}$.
@@ VII-C Traversability-Aware Navigation
In summary, the success of the real-world experiment validates that the robot learns to bind the color information with the level of traversability estimated from its own experiences, enabling it to avoid difficult-to-traverse regions based on visual input.
@@ VIII Conclusion and Future Work
In conclusion, we have introduced a multi-stage hierarchical reinforcement learning framework to achieve traversability-aware RGBD-based quadrupedal navigation. The framework is designed to train the navigation planner using real-world multi-modal data and the robot’s own experiences interacting with various terrains. The key design choices within the proposed hierarchical framework are thoroughly validated through extensive benchmarks. Our approach offers a novel way to train vision-based navigation planners that consider robot-centric traversability, making it easy to apply and customize planners for different robot platforms in complex outdoor environments.

### v2

@@ V-A4 Policy Representation and Training Details
The depth-based planner $\pi^{p}_{\text{depth}}$ is a deep neural network with two components: a depth image encoder $\phi_{\text{depth}}$ and a base MLP.
We train the depth-based planner $\pi^{p}_{\text{depth}}$ using PPO in Isaac Gym, where random obstacles (cylinders, boxes) and simulated noise are added to the depth images.
@@ V-C Online Learning for Traversability-Aware Navigation
After collecting the offline replay buffer, we begin training the final policy, the traversability-aware planner $\pi^{p}_{\text{rgbd}}$ that fuses depth and color images, and robot state feedback. This training is from scratch and is directly deployed on the robot hardware while exploring different terrains during navigation.
We retain the depth image encoder $\phi_{\text{depth}}$ and color image encoder $\phi_{\text{rgb}}$ and online train the base MLP. This approach avoids learning directly from raw image data, which is of much higher dimensionality, thus accelerating real-world training. The training objective is to enable the base MLP to properly utilize the latent embeddings from the depth and color images, along with the robot feedback, to avoid terrains with high traversability cost and obstacles while navigating to the goal.
The elements in the POMDP (such as observation, reward, and action) during this stage remain the same as those in the offline data collection stage, as described in Sec. V-B. The key difference is that we now start updating the planner policy (which produces the action $\mathbf{a}^{p}_{t}$) using the data collected with the new policy. We adopt DroQ algorithm [34] (modified SAC [28] with dropout in Q-value network), augmented it with RLPD. During each policy update, the sampled batch consists of two parts: half of the data is sampled from the offline dataset $\mathcal{D}_{\text{offline}}$ previously collected, and the other half is sampled from the online data $\mathcal{D}_{\text{online}}$ produced by the current policy. The update rule follows the standard SAC algorithm. By leveraging the offline dataset, we can accelerate online training and reduce the need for extensive exploration during online learning.
@@ VI-B Estimating Traversability
As shown in Fig. 5, our method, which includes the value estimate $T_{\text{value}}$ to evaluate overall control performance, performs significantly better than using the velocity tracking error alone ($T$ from Tracking), resulting in a lower total traversability cost over the trajectory (i.e. in sparse obstacles: $2.23\pm 0.52$ s) compared with $T$ from Tracking ($3.14\pm 0.66$ s), which uses the velocity tracking error alone.
The results indicate that relying solely on heuristic-based metrics, such as velocity tracking error, fails to account for other factors that may not affect tracking performance but cause other degradations, such as foot slippage or body instability. These factors can be considered by the proposed estimator based on the controller’s value function.
@@ VI-D Different Sources of Offline Data
However, if we only rely on the data collected by the human demonstrators, we also find there is performance degradation, especially for the goal-reaching objective (Ours $0.27\pm 0.06$ m/s versus Human Demo. Only $0.23\pm 0.03$ m/s). This is because human demonstrators may prefer the strategy that favors obstacle avoidance and lower traversability cost, but could be hard to consider minimizing traveling time.
The data collected by the depth-based planner can overcome this problem as it explicitly handles the goal vector information with a well-trained planning policy.
@@ VI-E Learning from Both Offline and Online Data
Compared to the Vanilla SAC method, which does not use any pre-collected dataset, we find that incorporating an offline replay buffer, as in our proposed method, significantly improves overall training performance. This advantage is particularly evident in our case, where we are addressing multi-modal learning from high-dimensional visual data. The pre-collected offline replay buffer, $\mathcal{D}_{\text{offline}}$, provides a high-quality dataset that helps to warm up the learning process. This not only enhances the sampling efficiency for online learning but also broadens the coverage of the distribution of collected trajectories, thereby improving the generalization of the learned planner.
We also compared our method with another approach that incorporates a pre-collected dataset during learning, but without using reinforcement learning: Behavior Cloning (BC). As shown in Fig. 8, BC demonstrates worse performance because it relies solely on regression from state-action demonstration pairs without any reward labels. This method is more susceptible to suboptimal demonstrations and fails to explore behaviors that could be more optimal than those demonstrated.
In contrast, our proposed method is less sensitive to the quality of the demonstrations because the transition pairs collected by our method include a reward term. After online exploration and learning, the agent optimizes for accumulated rewards and surpasses the performance of suboptimal demonstrations.
@@
VI-F Summary
In summary, we justify several key design choices for training the RGBD-based planner policy and provide a comprehensive quantitative analysis. The findings are summarized in four key points:
  • Traversability Estimator: We emphasize the use of value estimation $T_{\text{value}}$ from the locomotion controller. This approach results in a more reliable and robust robot-centric traversability estimate.
  • Types of Vision Input: We recommend using both color and depth information to enhance scene understanding and generalization of the outdoor environments, which benefits obstacle avoidance, terrain texture identification, and ultimately, traversability-aware navigation.
  • Sources of Offline Data: We underscore the importance of collecting offline data from diverse sources. By incorporating both human-centric and robot-centric demonstrations, the dataset becomes more varied and beneficial for subsequent online learning.
  • Learning from Offline and Online Data: We demonstrate that our proposed method, which integrates both offline and online data through reinforcement learning, improves sample efficiency and mitigates the effects of suboptimal demonstrations.
@@ VII-B Traversability Estimation in the Wild
Figure 10: The change of the uncertainty level when large disturbances that are not accurately simulated happen. The uncertainty level $\sigma$ of the estimated traversability $T_{\text{value}}$ represented by the orange line rises significantly after the disturbance occurs and then returns to a nominal level once the robot recovers from the disturbance. For comparison, we also plot the blue line representing the nominal walking situation on cobblestone terrain without disturbances, where the overall uncertainty level is lower.
@@
As shown in Fig. 9, the estimation results accurately reflect changes in terrain. The final result of the scalar traversability cost $T$ significantly increases when the robot traverses cobblestone areas and decreases on nominal concrete roads, correctly reflecting the difficulty of traversing through these regions. For example, cobblestone areas are challenging for the robot to traverse and introduce obvious oscillation on the robot base. The results of this estimatio is used as part of the reward $r^{p}_{t}$ in (3) to train the planner $\pi_{\text{rgbd}}^{p}$.
@@
VII-B1 Uncertainty-aware Estimation
To emphasize the necessity of explicit uncertainty evaluation for the estimation $T_{\text{value}}$ described in Sec. IV-B, we examine the uncertainty level of the traversability estimator under scenarios that are not accurately simulated, i.e. a large disturbance. When this large disturbance happens, the estimation from $V^{c}_{\text{terrain}}$ becomes uncertain about itself. As shown in Fig. 10, the uncertainty level $\sigma$ represented by the orange line rises significantly after the disturbance occurs and then returns to a nominal level once the robot recovers from the disturbance. We also plot the level of uncertainty during nominal walking without disturbance when the uncertainty level is acceptably low and the estimation $T_{\text{value}}$ is reliable, represented by blue lines. To have a reliable estimation, when the uncertainty level of $T_{\text{value}}$ rises, less confidence is placed on $T_{\text{value}}$ and the confidence on the backup estimation $T_{\text{track}}$ from velocity tracking errors increases as described in (1).
@@ VII-C Traversability-Aware Navigation
In summary, the results demonstrate that the robot-centric traversability estimation from the locomotion system provides reliable signals for the planner to generate trajectories with lower traversability costs. The success of the real-world experiment validates that the robot learns to bind the color information with the level of traversability estimated from its own experiences, enabling it to avoid difficult-to-traverse regions based on visual input.
@@ VIII Conclusion and Future Work
In conclusion, we have introduced a multi-stage hierarchical reinforcement learning framework to achieve traversability-aware RGBD-based quadrupedal navigation. The framework is designed to train the navigation planner using real-world multi-modal data and the robot’s own experiences interacting with various terrains.
The key design choices within the proposed hierarchical framework are thoroughly validated through extensive benchmarks.
Our approach offers a novel way to train vision-based navigation planners that consider robot-centric traversability, making it easy to apply and customize planners for different robot platforms in complex outdoor environments. However, one limitation of our approach is that it struggles to generalize to entirely new environments with different terrain textures.
For instance, a planner trained on Roads with Leaves might struggle with new terrain types, like a sand-filled bunker in Bunker Roads. To address this, a possible future direction could involve combining data collected from all different terrains into a unified offline dataset.
Thanks to the flexible online traversability estimator, we can easily expand the offline dataset by collecting transition pairs as the robot navigates diverse terrains during routine operations. This extensive offline dataset can be used for zero-shot transfer with offline learning methods or adapted quickly to new scenarios using our proposed online learning approach with minimal rollouts.